# A multi-messenger hierarchical triple merger gravitational-wave event pair GW190514-GW190521 inside AGN J124942.3 + 344929

### v2

@@ V Acknowledgments
We acknowledge valuable input from our anonymous referee which improved the original manuscript. This work is supported by National Key R&\&&D Program of China (2020YFC2201400), the National Natural Science Foundation of China (grant No. 11922303), and the Fundamental Research Funds for the Central Universities, Wuhan University (grant No. 2042022kf1182). We use the GW datasets released with GWTC-2.1 [2], labeled as “cosm” and “IMRPhenomXPHM.” This research has made use of data or software obtained from the Gravitational Wave Open Science Center (https://gwosc.org), a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan. This analysis was made possible following software packages: NumPy [62], SciPy [63], Matplotlib [64], emcee [65], IPython [66], corner [67], seaborn [68], and Astropy [69].

### v3

@@ V Acknowledgments
We acknowledge valuable input from our anonymous referee which improved the original manuscript. This work is supported by National Key R&\&&D Program of China (2020YFC2201400), the National Natural Science Foundation of China (grant No. 11922303), and the Fundamental Research Funds for the Central Universities, Wuhan University (grant No. 2042022kf1182). This research has made use of data or software obtained from the Gravitational Wave Open Science Center (https://gwosc.org), a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan.